OpenStudy (anonymous):

How to find the inverse of y=x^2-x by completing the square?

OpenStudy (campbell_st):

well can you complete the square in x...?

OpenStudy (anonymous):

I'm not sure.. I thought that we had to swap the x and y then solve for y then?

OpenStudy (campbell_st):

well the way the question reads... you need to complete the square 1st

OpenStudy (anonymous):

\[y\quad =\quad { x }^{ 2 }-x\]

OpenStudy (campbell_st):

so the 1st take is to halve the coefficient of x, which is -1 then square the answer so its \[(\frac{-1}{2})^2\] what would the value be..?

OpenStudy (anonymous):

i'm still confused on how to go on :/

OpenStudy (anonymous):

is it (x-1/2)^2=0?

OpenStudy (campbell_st):

ok... so if you have a quadratic \[x^2 + bx \] to complete the square you need to add \[(\frac{b}{2})^2\] to complete the square so in your question b = -1 so you need to add \[(\frac{-1}{2})^2\] to get the perfect square... does that make sense

OpenStudy (anonymous):

yep.. i get that

OpenStudy (campbell_st):

ok to balance the equation and keep it in its original form you need to also subtract that value so you have \[y = (x^2 -x + \frac{1}{4}) - \frac{1}{4}\] is that ok...?

OpenStudy (anonymous):

ahh, i see.. yep yep

OpenStudy (campbell_st):

ok.. so factoring the brackets you get \[y = (x - \frac{1}{2})^2 - \frac{1}{4}\] is that ok...?

OpenStudy (anonymous):

yeah :)

OpenStudy (campbell_st):

ok... now to find the inverse... swap x and y.. so you have \[x = (y - \frac{1}{2})^2 - \frac{1}{2}\] so the task to get the inverse is use the above equation and make x the subject.

OpenStudy (campbell_st):

oops it should read... to get the inverse... make y the subject

OpenStudy (anonymous):

and we solve for y now, yeah?

OpenStudy (campbell_st):

thats it... sorry for the typo

OpenStudy (anonymous):

it's all good

OpenStudy (campbell_st):

hope it helps you find the inverse

OpenStudy (anonymous):

yep! I got it! Thank you!!!